OpenStudy (anonymous):

Help ASAP HW 6×+11=5×-3

pooja195 (pooja195):

Are those x's?

OpenStudy (anonymous):

you solving for "x"?

OpenStudy (triciaal):

do not post a question and leave if you really need help.

OpenStudy (anonymous):

Yes I'm solving for "x"

Vocaloid (vocaloid):

ok, start by subtracting 5x from each side

OpenStudy (jennithemeani):

6x+11=5x-3 Subtract 5x on both sides x+11=-3 Isolate x by adding 11 on both sides x=14

Vocaloid (vocaloid):

nope, not quite 6x + 11 = 5x - 3 x + 11 = -3 x = -14

OpenStudy (jennithemeani):

Yep, sorry about that. I tend to forget the negatives

OpenStudy (anonymous):

How do you get -3

OpenStudy (jennithemeani):

From the original question the only difference from my answer and Vocaloid is that one is negative and one is positive, -14 is correct because -3-11=-14

Vocaloid (vocaloid):



OpenStudy (anonymous):

Oh ok I see now sorry about that it was a dumb answer

OpenStudy (anonymous):

Can you help me with another problem

OpenStudy (anonymous):

If that ok

OpenStudy (jennithemeani):

Fine by me :)

OpenStudy (jennithemeani):

7×-11=4×+19 subtract 4x from each side 3x-11=19 Add 11 to isolate x 3x=30 then divide by 3 x=10

OpenStudy (anonymous):

Thank you very much